# Language Models as Semantic Teachers: Post-Training Alignment for Medical Audio Understanding

### v1

@@ Appendix B Examples of Metadata–Report Pairs
To illustrate the supervision signal used during alignment, Table X provides representative examples of paired metadata and generated reports across all datasets. Each report is produced under strict schema constraints and is limited to restating clinician-verified fields without adding diagnostic speculation. These examples demonstrate how the LLM transforms structured metadata into concise, clinically grounded descriptions suitable for semantic alignment.
@@
TABLE X: Examples of clinician-verified metadata and the corresponding LLM-generated clinical reports across all datasets. Each generated report is a schema-constrained restatement of the metadata fields, ensuring factual fidelity and preventing the introduction of new diagnoses or speculative clinical reasoning.
@@ Appendix C OpenSmile Baseline Results
TABLE XI: Comparison of OpenSmile baseline with our approach. AUROC is reported for classification tasks (T1-T9, higher is better) and MAE for regression tasks (T10-T16, lower is better). We report mean and standard deviation from five independent runs for classification tasks and across subjects for regression tasks.
@@
Appendix D Qualitative Retrieval Examples
Appendix E Comparison of different LLMs
@@
TABLE XII: Comparison of different LLMs across respiratory and cardiac audio tasks. T1-T9: respiratory classification (AUROC), T10-T16: lung function estimation (MAE), T17-T18: cardiac classification (AUROC).
@@
Appendix F Performance of a respiratory-only model on respiratory and cardiac audio tasks
@@
TABLE XIII: Performance on respiratory and cardiac audio tasks using model trained exclusively on respiratory sounds data. T1-T9: respiratory classification (AUROC), T10-T16: lung function estimation (MAE), T17-T18: cardiac classification (AUROC).
@@
Appendix G Respiratory and cardiac audio performance without augmentation
@@
TABLE XIV: Performance on respiratory and cardiac audio tasks without data augmentation during alignment. T1-T9: respiratory classification (AUROC), T10-T16: lung function estimation (MAE), T17-T18: cardiac classification (AUROC).
@@
Appendix H Alignment Strategy Analysis
@@
TABLE XV: Single–layer vs. multi–layer alignment on each downstream task. Last-L = alignment applied only to the final transformer block; Multi-L = alignment applied to blocks {3, 6, 9, last (12)}, with the four alignment losses averaged. T1-T9: respiratory classification (AUROC), T10-T16: lung function estimation (MAE), T17-T18: cardiac classification (AUROC).
@@
T9 COPD severity (Lung sounds)  [^1] Multi-class AUROC is macro-averaged. 0.710 0.744

### v2

@@ Appendix B Examples of Metadata–Report Pairs
To illustrate the supervision signal used during alignment, Table XII provides representative examples of paired metadata and generated reports across all datasets. Each report is produced under strict schema constraints and is limited to restating clinician-verified fields without adding diagnostic speculation. These examples demonstrate how the LLM transforms structured metadata into concise, clinically grounded descriptions suitable for semantic alignment.
@@
TABLE XII: Examples of clinician-verified metadata and the corresponding LLM-generated clinical reports across all datasets. Each generated report is a schema-constrained restatement of the metadata fields, ensuring factual fidelity and preventing the introduction of new diagnoses or speculative clinical reasoning.
@@ Appendix C OpenSmile Baseline Results
TABLE XIII: Comparison of OpenSmile baseline with our approach. AUROC is reported for classification tasks (T1-T9, higher is better) and MAE for regression tasks (T10-T16, lower is better). We report mean and standard deviation from five independent runs for classification tasks and across subjects for regression tasks.
@@
Appendix D Comparison of different LLMs
@@
TABLE XIV: Comparison of different LLMs across respiratory and cardiac audio tasks. T1-T9: respiratory classification (AUROC), T10-T16: lung function estimation (MAE), T17-T18: cardiac classification (AUROC).
@@
Appendix E Label-Masked Report Ablation
This appendix evaluates potential label leakage from LLM-generated reports. We regenerate reports with label-revealing fields/keywords (e.g., diagnosis/outcome terms) masked, re-train AcuLa using audio + masked reports, and compare downstream performance against AcuLa trained with the original reports across all tasks (T1–T18). Table XV shows representative report snippets before and after masking. As shown in Table XVI, the overlap tasks exhibit a slightly larger but still minor AUROC decrease (mean $\Delta\mathrm{AUROC}=-0.0047$, with a maximum drop of 0.009 on T3), whereas the non-overlap classification tasks show almost no change (mean $\Delta\mathrm{AUROC}=-0.0010$). Regression tasks (T10–T16) remain essentially unchanged, with differences of at most 0.007. Overall, the small and localized degradation on overlap tasks suggests that explicit diagnosis/outcome tokens contribute minimally to alignment, and the largely stable performance across all tasks indicates that AcuLa’s gains are not primarily driven by trivial label-text leakage.
TABLE XV: Examples of report text before and after label masking. Diagnosis/outcome keywords are removed to mitigate label leakage, while symptoms and descriptive findings are retained.
Original (before masking) Masked (after masking)
Cardiac auscultation findings are within normal limits, indicating no abnormalities in heart sounds despite the presence of non-severe pneumonia. The record annotation indicates a normal assessment with an event type classified as normal. The findings are within expected parameters with no anomalies noted.
Auscultation reveals continuous adventitious sounds characterized as wheezes. These findings are consistent with a diagnosis of non-severe pneumonia. The patient’s lung auscultation recordings demonstrate the presence of both continuous and discontinuous adventitious sounds.
The patient is a middle-aged female with a positive COVID-19 test presenting with runny or blocked nose, new continuous cough, fatigue, and headache. Respiratory auscultation may reveal mild bronchial breath sounds or scattered crackles related to viral upper respiratory involvement, consistent with her symptom profile. No signs of respiratory distress or wheezing noted. This participant is a middle-aged female who does not wear a mask and has never smoked. She reported experiencing a cough, new continuous cough, runny or blocked nose, diarrhea, fatigue, and headache.
Anterior left chest auscultation reveals no crackles or wheezes in a 3-year-old male diagnosed with bronchiolitis. The patient is a 3-year-old male with findings from the anterior left chest. The acquisition mode was sequential/single channel. No crackles or wheezes were noted.
TABLE XVI: Label-masked report ablation. We mask label-revealing fields/words during report generation (e.g., diagnosis/outcome terms) and re-train AcuLa with audio + masked reports.
ID Task AcuLa (original reports) AcuLa (masked reports)
T1 Covid (Exhale) 0.698 0.695
T2 Covid (Cough) 0.730 0.724
T3 Covid (Cough) 0.887 0.878
T4 Gender (Cough) 0.796 0.793
T5 COPD (Lung) 0.826 0.819
T6 Smoker (Cough) 0.830 0.830
T7 Gender (Cough) 0.845 0.845
T8 Obstructive (Lung) 0.752 0.750
T9 COPD severity (Lung) 0.710 0.708
T10 FVC (Breath) 0.865 0.865
T11 FEV1 (Breath) 0.742 0.744
T12 FEV1/FVC (Breath) 0.127 0.127
T13 FVC (Vowel) 0.779 0.781
T14 FEV1 (Vowel) 0.725 0.724
T15 FEV1/FVC (Vowel) 0.123 0.123
T16 Breathing Rate 2.388 2.395
T17 Murmur / Healthy 0.675 0.673
T18 Symptomatic / Healthy 0.647 0.644
Appendix F Baseline Adaptation Results
This appendix reports per-task performance under a fair baseline adaptation setting. We fine-tune AudioMAE and CLAP in-domain on the same alignment training audio (train splits only) without using reports, and evaluate all models using the identical frozen linear-probe protocol. Audio-only adaptation improves both encoders compared to their pretrained counterparts, but AcuLa remains consistently stronger on most tasks, particularly for respiratory and cardiac condition inference. These results support that the proposed semantic alignment provides benefits beyond in-domain audio exposure alone. Tables XVIIXIX provide per-task results for respiratory condition inference (T1–T9, AUROC), lung function estimation (T10–T16, MAE), and cardiac condition inference (T17–T18, AUROC).
TABLE XVII: Per-task AUROC ($\uparrow$) on respiratory health condition inference (T1–T9) under fair baseline adaptation. AudioMAE-ft and CLAP-ft are in-domain fine-tuned on the same alignment training audio (train splits only) without reports.
ID Task AudioMAE-ft CLAP-ft AcuLa (Ours)
T1 Covid (Exhale) 0.608 0.612 0.698
T2 Covid (Cough) 0.678 0.683 0.730
T3 Covid (Cough) 0.671 0.679 0.887
T4 Gender (Cough) 0.694 0.723 0.796
T5 COPD (Lung) 0.885 0.932 0.826
T6 Smoker (Cough) 0.612 0.732 0.830
T7 Gender (Cough) 0.755 0.779 0.845
T8 Obstructive (Lung) 0.704 0.726 0.752
T9 COPD severity (Lung) 0.570 0.640 0.710
Avg. Average 0.686 0.723 0.786
TABLE XVIII: Per-task MAE ($\downarrow$) on lung function estimation (T10–T16) under fair baseline adaptation. AudioMAE-ft and CLAP-ft are in-domain fine-tuned on the same alignment training audio (train splits only) without reports.
ID Task AudioMAE-ft CLAP-ft AcuLa (Ours)
T10 FVC (Breath) 0.878 0.879 0.865
T11 FEV1 (Breath) 0.780 0.796 0.742
T12 FEV1/FVC (Breath) 0.128 0.132 0.127
T13 FVC (Vowel) 0.845 0.856 0.779
T14 FEV1 (Vowel) 0.808 0.795 0.725
T15 FEV1/FVC (Vowel) 0.129 0.133 0.123
T16 Breathing Rate 2.477 2.506 2.388
Avg. Average 0.864 0.871 0.821
TABLE XIX: Per-task AUROC ($\uparrow$) on cardiac condition inference (T17–T18) under fair baseline adaptation. AudioMAE-ft and CLAP-ft are in-domain fine-tuned on the same alignment training audio (train splits only) without reports.
ID Task AudioMAE-ft CLAP-ft AcuLa (Ours)
T17 Murmur / Healthy 0.640 0.652 0.675
T18 Symptomatic / Healthy 0.618 0.622 0.647
Avg. Average 0.629 0.637 0.661
Appendix G Qualitative Retrieval Examples
Appendix H Semantic Alignment Evaluation via Audio–Text Retrieval
[…]
To reduce trivial matching driven by explicit label words, we additionally evaluate retrieval using label-masked reports, where fields that directly reveal downstream labels (e.g., diagnosis/outcome keywords such as “COVID-19 positive”, “pneumonia”, “abnormal”) are removed. Masking only alters the report text; the underlying audio–text pairing remains unchanged.
[…]
Setting R@1 $\uparrow$ R@5 $\uparrow$ R@10 $\uparrow$ MedR $\downarrow$
[…]
@@ Appendix H Semantic Alignment Evaluation via Audio–Text Retrieval
Appendix I Performance of a respiratory-only model on respiratory and cardiac audio tasks
@@
TABLE XXI: Performance on respiratory and cardiac audio tasks using model trained exclusively on respiratory sounds data. T1-T9: respiratory classification (AUROC), T10-T16: lung function estimation (MAE), T17-T18: cardiac classification (AUROC).
@@
Appendix J Respiratory and cardiac audio performance without augmentation
@@
TABLE XXII: Performance on respiratory and cardiac audio tasks without data augmentation during alignment. T1-T9: respiratory classification (AUROC), T10-T16: lung function estimation (MAE), T17-T18: cardiac classification (AUROC).
@@
Appendix K Alignment Strategy Analysis
@@
TABLE XXIII: Single–layer vs. multi–layer alignment on each downstream task. Last-L = alignment applied only to the final transformer block; Multi-L = alignment applied to blocks {3, 6, 9, last (12)}, with the four alignment losses averaged. T1-T9: respiratory classification (AUROC), T10-T16: lung function estimation (MAE), T17-T18: cardiac classification (AUROC).
@@
T9 COPD severity (Lung sounds) 0.710 0.744